Please help me solve this question
Let the space of the random variable X be A = {x : 0 < x < 10} and let P(A1) =3 /8, where A1= {X : 1< X < 5}. Show that P(A2) ≤ 5/8. Where A2 = {x:5≤ x <10}

Answer by stanbon(75887)

Please help me solve this question
Let the space of the random variable X be A = {x : 0 < x < 10}
and let P(A1) = 3 /8, where A1= {X : 0< X < 5}. Show that P(A2) ≤ 5/8. Where A2 = {x:5≤ x <10
-------
A covers the interval 0 A = A1+A2
----
0< x < 5 U 5<= x < 10 = 0 -----
A1 U A2 = A
--
P(A1)+P(A2) = 1
----
3/8 + P(A2) = 1
---
Therefore: P(A2) = 1-(3/8) = 5/8
===============================
Cheers,
Stan H.
-------------------------------